A 12 ounce drink costs $0.75 .Calculate the following quantities or cost for this drink

QUARTS,GALLONS LITERS MILLILITERS ,TABLESPOONS, POUNDS ,TON,KILOGRAMS MILIGRAMS, DOLLARS/GALLONS, DOLLARS/LITER
Answer #1

Answer)

- We know that one quart = 32 ounce

And given price for 12 ounce = $0.75

12 ounce = $0.75

1 ounce = 0.75/12

Multipling both sides with 32, we get

32 ounce = (0.75/12)*32

32 ounce = $2

So, 1 quart costs $2

- 1 gallon = 128 ounces

12 ounces = $0.75

1 ounce = (0.75/12)

128 ounces = $(0.75/12)*128 = $8

So, one gallon costs $8

-

1 pound = 16 ounces.

12 ounces = $0.75

16 ounces = (0.75/12)*16

Therefore, one pound costs $1

-

One ton = 32000 ounces

12 ounce = $0.75

32000 ounces = (0.75/12)*32000

Therefore,.

One ton costs $2000

-

One kilogram = 35.274 ounces

12 ounce = $0.75

35.274 = (0.75/12)*(35.274)

Therefore, one kilogram costs = $2.204625

-

1 tablespoon = 0.5 ounce

12 ounce = $0.75

0.5 = (0.75/12)*0.5

Therefore, one tablespoon costs $0.03125

-

1 milligram = 3.527 × 10-5 ounce

Therefore,

Cost = $(0.75/12)*3.527*(10^-5)

= $0.000002204375